A mole of a newly discovered crystalline substance weighs 14.56 grams and displaces 2.3 ml of water. Find the atomic radius and density of each atom if it is a simple cubic cell? BCC? FCC? HCP?


1
Expert's answer
2022-03-30T02:31:01-0400

V=2.5mL=2.5cm=2.5×10-6 m3

V=L3

2.5×10-6m3 =L3

L=0.09259×10-2m

L=9.259×10-4m


"\therefore" Atomicradius="\frac{1}{2}×length"

"\frac{1}{2}" ×9.259×10-4

=0.00046295m

=462.95μm

Mass=14.56g=0.01456kg

Volume =2.5×10-6 m3

"\therefore" "\rho=\frac{mass}{volume}"


"=0.01456kg\div" 2.5×10-6 m3


=5.824×103kg/m3